# Learning finitely correlated states: stability of the spectral reconstruction

### v2

@@ 4.2 Error propagation bound
Theorem 4.3 (Error propagation).
Let $\omega$ be a finitely correlated state. For the state reconstruction in Definition 4.1, and the error parameters defined in Eqs. (77), (78) and (79), if $\hat{U}^{\intercal}U$ is invertible, then
@@
$\leq\Delta\|\hat{M}^{-1}\tilde{e}_{t-1}\|_{\infty\rightarrow\infty,cb}=\Delta\,,$ (90)
@@
$\leq\|\hat{M}^{-1}(\tilde{\mathbb{K}}-\hat{\mathbb{K}})\hat{M}\|_{\infty\rightarrow e^{\prime},cb}\|\hat{M}^{-1}(\hat{e}_{t-1}-\tilde{e}_{t-1})\|_{\infty\rightarrow e^{\prime},cb}$
@@
4.3 General bound
With the notation of Section 3, let us consider the order norm on $\mathbb{C}^{m}$ denoted as $\lVert\cdot\rVert_{e}$. Then the error propagation bound from Theorem 4.3, with $M$ being the identity map.
Lemma 4.4.
We have that
@@ 4.3 General bound
where we used Lemma A.2 for the first inequatity, (51) for the second inequality, and (59) for the last one. Lastly, for $\delta_{1}$, from (77)
@@
Theorem 4.5 (Reconstruction error from local estimates).
@@
$\|\tau(\Omega-\hat{\Omega})\|_{2}$ $\leq\frac{3\sigma_{m}(\Omega)}{4}\,\epsilon\,,$ (112)
$\|(\Omega-\hat{\Omega})(\mathds{1})\|_{2}$ $\leq\frac{\sqrt{3}\sigma_{m}(\Omega)}{2}\epsilon\,,$ (113)
@@
It is immediate to verify that the above inequalities are such that $\|\Omega-\hat{\Omega}\|_{2\rightarrow 2}\leq\|\Omega-\hat{\Omega}\|_{2}\leq\frac{\sigma_{m}(\Omega)^{2}}{3}\leq\frac{\sigma_{m}(\Omega)}{3}$, because $\sigma_{m}(\Omega)\leq\|\Omega\|_{2}\leq 1$. We can thus insert the bounds from Proposition 4.4 in the bound of Theorem 4.3, which now reads $\|\hat{\omega}_{t}-\omega_{t}\|_{1}\leq(1+\epsilon)(1+\frac{4}{3}\epsilon)(1+\frac{4}{3t}\epsilon)^{t}-1,$ and use $(1+a/t)^{t}\leq 1+2a$ for $0\leq a\leq 1$. ∎
Thanks to these bounds, for general finitely-correlated states, we can prove the following
Theorem 4.6.
Let $\omega$ be a finitely correlated state, with associated map $\Omega$. There exists an algorithm which takes as input $n$ copies of $\omega$, an integer $s$ and a threshold $\eta$ such that it makes measurements on marginals of size $s$ and it outputs an estimate $\hat{\omega}_{t}$ such that, with high probability
$d_{\mathrm{Tr}}(\hat{\omega}_{t},\omega_{t})\leq\epsilon$ (117)
whenever $s\geq\max\{t_{2}^{*},t_{1}^{*}+1\}$, $\sigma_{m}(\Omega)/2\geq\eta\geq 2\varepsilon$ with $\varepsilon=\frac{\epsilon\sigma^{3}_{m}(\Omega)}{20tm\sqrt{d_{\mathcal{A}}}}$ and $n$ larger than the number of copies of $\omega_{2s}$ and $\omega_{2s+1}$ necessary for learning $\omega_{2s},\omega_{2s+1}$ in Hilbert-Schmidt distance at precision $\varepsilon$.
@@
According to the error propagation bound in Theorem 4.5, the bottleneck is the second term in Eq. (114), which is $O\left(\frac{\epsilon\sigma_{m}(\Omega)^{3}}{tm\sqrt{d_{\mathcal{A}}}}\right)$. All the relevant marginals can be estimated at that precision with high probability with any preferred method. Then, the first $m$ singular values of the estimated marginals will be closer than $O\left(\frac{\epsilon\sigma_{m}(\Omega)^{3}}{tm\sqrt{d_{\mathcal{A}}}}\right)$ to the true values, via Lemma B.1, meaning that singular value decomposition truncated to singular values larger than $\sigma_{m}(\Omega)+O\left(\frac{\epsilon\sigma_{m}(\Omega)^{3}}{tm\sqrt{d_{\mathcal{A}}}}\right)$ selects a map $\hat{U}$ of rank $m$ and the error propagation bound applies. ∎
4.4 Bound for $C^{*}$-FCS
If the state admits a $C^{*}$-realization, we can make use of the operator system inherited by the usual $C^{*}$ matrix order to get an alternative error propagation bound, which can be better than the one obtained for general states. In fact, the new sample complexity bound is obtained by exchanging $m$ with $d_{\mathcal{B}}$, and we have $m\leq d^{2}_{\mathcal{B}}$, making the particular bound for $C^{*}$ states non trivial. The most relevant difference in the proof is that we can explicitly compute the cb norm of a map into $\mathcal{B}$ by considering a single amplification which is determined by $d_{\mathcal{B}}$, see Lemma A.3. With the notation of Section 3, let us consider the order norm on $\mathcal{V}$ denoted as $\lVert\cdot\rVert_{e^{\prime}}$, $e^{\prime}=L\mathds{1}_{\mathcal{B}}$. Then the error propagation bound from Theorem 4.3 applies, where we can choose the map $M$ to be $M=U^{\intercal}\mathcal{F}R$.
Lemma 4.7.
@@ 4.4 Bound for $C^{*}$-FCS
$\Delta$ $\leq\frac{8d_{\mathcal{B}}\sqrt{d_{\mathcal{A}}}}{\sqrt{3}\sigma_{m}(\Omega)}\left(\frac{\|\Omega-\hat{\Omega}\|_{2}}{\sigma_{m}({\Omega})^{2}}+\frac{\|\Omega_{(\cdot)}-\hat{\Omega}_{(\cdot)}\|_{2}}{3\sigma_{m}(\Omega)}\right),$ (119)
@@
We will use the Euclidean norm on $\mathcal{V}$ induced by the identification of $\mathcal{V}$ with a subspace $\mathcal{V}^{\prime}$ of $\mathcal{W}$ via the map $R$. For $x\in\mathcal{V}$, we thus have $\lVert x\rVert_{e^{\prime}}=\lVert Rx\rVert_{Re^{\prime}}\leq\lVert Rx\rVert_{2}=\lVert x\rVert_{2}$. This implies $\lVert M^{-1}\rVert_{2\rightarrow e^{\prime}}\leq\lVert M^{-1}\rVert_{2\rightarrow 2}$. Note that the singular values of $U^{\intercal}\mathcal{F}R$ and $U^{\intercal}\mathcal{F}\Pi_{\mathcal{V}^{\prime}}$ are the same. On the other hand, we can observe that the unnormalized $l$-th right singular vector of $U^{\intercal}\mathcal{F}\Pi_{\mathcal{V}^{\prime}}$ can be written as $\mathcal{E}^{t^{*}}_{X_{l}}(\mathds{1}_{\mathcal{B}})$ for some $X_{l}\in\mathcal{A}_{[1,t^{*}]}$ and orthogonal to the kernel of $\Omega$, and $\|X_{l}\|_{2}=1$. But then
@@
$\lVert M^{-1}\rVert_{2\rightarrow e^{\prime}}\leq\lVert M^{-1}\rVert_{2\rightarrow 2}\leq\frac{\sqrt{d_{B}}}{\sigma_{m}(\Omega)}.$ (122)
In the following, we repeat the steps of Lemma 4.4, explaining only what is different in this case. We have
@@
$=\|\mathrm{id}_{\mathcal{B}}\otimes RM^{-1}(\hat{U}^{\intercal}U)^{-1}(\tilde{\mathbb{K}}-\hat{\mathbb{K}})\hat{U}^{\intercal}UM\|_{\mathds{1}_{\mathcal{A}}\otimes\mathds{1}_{\mathcal{B}}\otimes e^{\prime}\rightarrow\mathds{1}_{\mathcal{B}}\otimes\mathds{1}_{\mathcal{B}}}$ (129)
$\leq\|\mathrm{id}_{\mathcal{B}}\otimes RM^{-1}(\hat{U}^{\intercal}U)^{-1}(\tilde{\mathbb{K}}-\hat{\mathbb{K}})\hat{U}^{\intercal}UM\|_{\mathds{1}_{\mathcal{A}}\otimes\mathds{1}_{\mathcal{B}}\otimes e^{\prime}\rightarrow\mathds{1}_{\mathcal{B}}\otimes\mathds{1}_{\mathcal{B}}}$ (130)
$\leq\lVert\mathrm{id}_{\mathcal{B}}\otimes RM^{-1}\rVert_{2\rightarrow\mathds{1}_{B}\otimes\mathds{1}_{B}}\|(\hat{U}^{\intercal}U)^{-1}\|_{2\rightarrow 2}\|\tilde{\mathbb{K}}-\hat{\mathbb{K}}\|_{2\rightarrow 2}\|\hat{U}^{\intercal}U\|_{2\rightarrow 2}\|\mathrm{id}_{\mathcal{A}}\otimes\mathrm{id}_{\mathcal{B}}\otimes M\|_{\mathds{1}_{\mathcal{A}}\otimes e^{\prime}\rightarrow 2,2}$ (131)
@@
where we used Lemma A.3 in the third equality and (122) and (52) in the third inequality. Lastly, for $\delta_{1}$:
@@
$\leq\|(\hat{\rho}-\tilde{\rho})\hat{U}^{\intercal}UM\|_{2}$ (135)
$\leq\|\hat{\rho}-\tilde{\rho}\|_{2}\,,$ (136)
@@
Theorem 4.8 (Reconstruction error from local estimates).
@@
$\|\tau(\Omega-\hat{\Omega})\|_{2}$ $\leq\frac{3\sigma_{m}(\Omega)}{4}\,\epsilon\,,$ (138)
$\|(\Omega-\hat{\Omega})(\mathds{1})\|_{2}$ $\leq\frac{\sqrt{3}\sigma_{m}(\Omega)}{2\sqrt{d_{B}}}\epsilon\,,$ (139)
@@
It is immediate to verify that the above inequalities are such that $\|\Omega-\hat{\Omega}\|_{2\rightarrow 2}\leq\|\Omega-\hat{\Omega}\|_{2}\leq\frac{\sigma_{m}(\Omega)^{2}}{3}\leq\frac{\sigma_{m}(\Omega)}{3}$, because $\sigma_{m}(\Omega)\leq\|\Omega\|_{2}\leq 1$. We can thus insert the bounds from Proposition 4.4 in the bound of Theorem 4.3, and use $(1+a/t)^{t}\leq 1+2a$ for $0\leq a\leq 1$. ∎
The proof of the main theorem is analogous to the one of Theorem 4.6. Note that the only difference is the dependence of $\varepsilon$ on the parameters of the model.
Theorem 4.9.
Let $\omega$ be a $C^{*}$-finitely correlated state with memory system $\mathbb{M}_{d_{\mathcal{B}}}$ and associated map $\Omega$. There exists an algorithm which takes as input $n$ copies of $\omega$, an integer $s$ and a threshold $\eta$ such that it makes measurements on marginals of size $s$ and it outputs an estimate $\hat{\omega}_{t}$ (not necessarily positive-semidefinite) such that, with high probability
$d_{\mathrm{Tr}}(\hat{\omega}_{t},\omega_{t})\leq\epsilon$ (143)
whenever $s\geq\max\{t_{2}^{*},t_{1}^{*}+1\}$, $\sigma_{m}(\Omega)/2\geq\eta\geq 2\varepsilon$ with $\varepsilon=\frac{\epsilon\sigma^{3}_{m}(\Omega)}{20td_{B}\sqrt{d_{\mathcal{A}}}}$, and $n$ larger than the number of copies of $\omega_{2s}$ and $\omega_{2s+1}$ necessary for learning $\omega_{2s},\omega_{2s+1}$ in Hilbert-Schmidt distance at precision $\varepsilon$.
5 Non-homogeneous FCS on a finite chain

### v3

@@ C.2 Error propagation bound
Theorem C.3 (Error propagation).
Let $\omega$ be a finitely correlated state. For the state reconstruction in Definition C.1, and the error parameters defined in Eqs. (109), (110) and (111), if $\hat{U}^{\intercal}U$ is invertible, then
@@
$\leq\Delta\|\hat{M}^{-1}\tilde{e}_{t-1}\|_{\infty\rightarrow e^{\prime},cb}=\Delta\,,$ (122)
@@
$\leq\|\hat{M}^{-1}(\tilde{\mathbb{K}}-\hat{\mathbb{K}})\hat{M}\|_{\mathds{1}\otimes e^{\prime}\rightarrow e^{\prime},cb}\|\hat{M}^{-1}(\hat{e}_{t-1}-\tilde{e}_{t-1})\|_{\infty\rightarrow e^{\prime},cb}$
@@
C.3 General bound
With the notation of Section B, let us consider the order norm on $\mathbb{C}^{m}$ denoted as $\lVert\cdot\rVert_{e}$. Then the error propagation bound follows from Theorem C.3, with $M$ being the identity map.
Lemma C.4.
If $\|\Omega-\hat{\Omega}\|_{2\rightarrow 2}\leq\frac{\sigma_{m}(\Omega)}{3}$, we have
@@ C.3 General bound
where we used Lemma B.2 for the first inequatity, (82) for the second inequality, and (91) for the last one. Lastly, for $\delta_{1}$, from (109)
@@
Theorem C.5 (Reconstruction error from local estimates).
@@
$\|\tau\Omega-\widehat{\tau\Omega})\|_{2}$ $\leq\frac{3\sigma_{m}(\Omega)}{4}\,\epsilon\,,$ (144)
$\|\Omega(\mathds{1})-\widehat{\Omega(\mathds{1})}\|_{2}$ $\leq\frac{\sqrt{3}\sigma_{m}(\Omega)}{2}\epsilon\,,$ (145)
@@
It is immediate to verify that the above inequalities are such that $\|\Omega-\hat{\Omega}\|_{2\rightarrow 2}\leq\|\Omega-\hat{\Omega}\|_{2}\leq\frac{\sigma_{m}(\Omega)^{2}}{3}\leq\frac{\sigma_{m}(\Omega)}{3}$, because $\sigma_{m}(\Omega)\leq\|\Omega\|_{2}\leq 1$. We can thus insert the bounds from Proposition C.4 in the bound of Theorem C.3, which now reads $\|\hat{\omega}_{t}-\omega_{t}\|_{1}\leq(1+\epsilon)(1+\frac{4}{3}\epsilon)(1+\frac{4}{3t}\epsilon)^{t}-1,$ and use $(1+a/t)^{t}\leq 1+2a$ for $0\leq a\leq 1$. ∎
Thanks to these bounds, for general finitely-correlated states, we can prove (see Definition 2.1 for the definiton of the class $\mathcal{S}(m,s,\eta)$):
Theorem C.6 (Theorem 2.6, restated).
Let $\tilde{\omega}\in\mathcal{S}(m,s,\eta)$. Let $\hat{\omega}_{s},\hat{\omega}_{2s},\hat{\omega}_{2s+1}$ be estimates of ${\omega}_{s},{\omega}_{2s},{\omega}_{2s+1}$ respectively, such that $D_{HS}(\hat{\omega}_{s},{\omega}_{s})$, $D_{HS}(\hat{\omega}_{2s},{\omega}_{2s})$, $D_{HS}(\hat{\omega}_{2s+1},{\omega}_{2s+1})$ are smaller than $\frac{\epsilon\eta^{3}}{20tm\sqrt{d_{\mathcal{A}}}}$. Then, $\hat{\omega}_{t}$ constructed from the output of Algorithm 1 satisfies
$\frac{\|\hat{\omega}_{t}-\omega_{t}\|_{1}}{2}\leq\epsilon.$ (149)
@@
According to the error propagation bound in Theorem C.5, the bottleneck is the second term in Eq. (146), which is $O\left(\frac{\epsilon\sigma_{m}(\Omega)^{3}}{tm\sqrt{d_{\mathcal{A}}}}\right)$. All the relevant marginals can be estimated at that precision with high probability with any preferred method. Then, the first $m$ singular values of the estimated marginals will be closer than $O\left(\frac{\epsilon\sigma_{m}(\Omega)^{3}}{tm\sqrt{d_{\mathcal{A}}}}\right)$ to the true values, via Lemma G.1, meaning that singular value decomposition truncated to singular values larger than $\sigma_{m}(\Omega)-O\left(\frac{\epsilon\sigma_{m}(\Omega)^{3}}{tm\sqrt{d_{\mathcal{A}}}}\right)$ selects a map $\hat{U}$ of rank $m$ and the error propagation bound applies. ∎
C.4 Bound for $C^{*}$-FCS
If the state admits a $C^{*}$-realization, we can make use of the operator system inherited by the usual $C^{*}$ matrix order to get an alternative error propagation bound, which can be better than the one obtained for general states. In fact, the new sample complexity bound is obtained by exchanging $m$ with $d_{\mathcal{B}}$, and we have $m\leq d^{2}_{\mathcal{B}}$, making the particular bound for $C^{*}$ states non trivial. The most relevant difference in the proof is that we can explicitly compute the cb norm of a map into $\mathcal{B}$ by considering a single amplification which is determined by $d_{\mathcal{B}}$, see Lemma B.3. With the notation of Section B, let us consider the order norm on $\mathcal{V}$ denoted as $\lVert\cdot\rVert_{e^{\prime}}$, $e^{\prime}=L\mathds{1}_{\mathcal{B}}$. Then the error propagation bound from Theorem C.3 applies, where we can choose the map $M$ to be $M=U^{\intercal}\mathcal{F}R$.
Lemma C.7.
@@ C.4 Bound for $C^{*}$-FCS
$\Delta$ $\leq\frac{8d_{\mathcal{B}}\sqrt{d_{\mathcal{A}}}}{\sqrt{3}\sigma_{m}(\Omega)}\left(\frac{\|\Omega-\hat{\Omega}\|_{2}}{\sigma_{m}({\Omega})^{2}}+\frac{\|\Omega_{(\cdot)}-\hat{\Omega}_{(\cdot)}\|_{2}}{3\sigma_{m}(\Omega)}\right)$ (151)
@@
We will use the Euclidean norm on $\mathcal{V}$ induced by the identification of $\mathcal{V}$ with a subspace $\mathcal{V}^{\prime}$ of $\mathcal{W}$ via the map $R$. Via this identification we can also consider the Schatten norms of $x\in\mathcal{V}$. For $x\in\mathcal{V}$, we thus have $\lVert x\rVert_{\infty}=\lVert Rx\rVert_{\infty}\leq\lVert Rx\rVert_{2}=\lVert x\rVert_{2}$. This implies $\lVert RM^{-1}\rVert_{2\rightarrow\infty}\leq\lVert RM^{-1}\rVert_{2\rightarrow 2}$. We also have $\|x\|_{(Re^{\prime})^{(n)}}\leq\|x\|_{\infty}$ for $x\in\mathbb{M}_{n}(\mathcal{V})$ as an immediate consequence of the definition of the order norm on $\mathcal{V}$. Note that the singular values of $U^{\intercal}\mathcal{F}R$ and $U^{\intercal}\mathcal{F}\Pi_{\mathcal{V}^{\prime}}$ are the same. On the other hand, we can observe that the unnormalized $l$-th right singular vector of $U^{\intercal}\mathcal{F}\Pi_{\mathcal{V}^{\prime}}$ can be written as $\mathcal{E}^{t^{*}}_{X_{l}}(\mathds{1}_{\mathcal{B}})$ for some $X_{l}\in\mathcal{A}_{[1,t^{*}]}$ and orthogonal to the kernel of $\Omega$, and $\|X_{l}\|_{2}=1$. But then
@@
$\lVert M^{-1}\rVert_{2\rightarrow e^{\prime}}=\lVert RM^{-1}\rVert_{2\rightarrow Re^{\prime}}\leq\lVert RM^{-1}\rVert_{2\rightarrow\infty}\leq\lVert RM^{-1}\rVert_{2\rightarrow 2}\leq\frac{\sqrt{d_{\mathcal{B}}}}{\sigma_{m}(\Omega)}.$ (154)
In the following, we repeat the steps of Lemma C.4, explaining only what is different in this case. We have
@@
$\leq\|RM^{-1}(\hat{U}^{\intercal}U)^{-1}(\tilde{\mathbb{K}}-\hat{\mathbb{K}})\hat{U}^{\intercal}UM\|_{\mathds{1}_{\mathcal{A}}\otimes e^{\prime}\rightarrow\infty,cb}$ (161)
$=\|\mathrm{id}_{\mathcal{B}}\otimes RM^{-1}(\hat{U}^{\intercal}U)^{-1}(\tilde{\mathbb{K}}-\hat{\mathbb{K}})\hat{U}^{\intercal}UM\|_{\mathds{1}_{\mathcal{A}}\otimes\mathds{1}_{\mathcal{B}}\otimes e^{\prime}\rightarrow\infty}$ (162)
$\leq\lVert\mathrm{id}_{\mathcal{B}}\otimes RM^{-1}\rVert_{2\rightarrow\infty}\|(\hat{U}^{\intercal}U)^{-1}\|_{2\rightarrow 2}\|\tilde{\mathbb{K}}-\hat{\mathbb{K}}\|_{2\rightarrow 2}\|\hat{U}^{\intercal}U\|_{2\rightarrow 2}\|\mathrm{id}_{\mathcal{A}}\otimes\mathrm{id}_{\mathcal{B}}\otimes M\|_{\mathds{1}_{\mathcal{A}}\otimes e^{\prime}\rightarrow 2}$ (163)
@@
where we used that $\|x\|_{(Re^{\prime})^{(n)}}\leq\|x\|_{\infty}$ in the first inequality, Lemma B.3 in the third equality and (154) and (84) in the third inequality. Lastly, for $\delta_{1}$:
@@
$\leq\|(\hat{\rho}-\tilde{\rho})\hat{U}^{\intercal}U\|_{2}$ (167)
$\leq\|\hat{\rho}-\tilde{\rho}\|_{2}\,$ (168)
@@
Theorem C.8 (Reconstruction error from local estimates).
@@
$\|\tau\Omega-\widehat{\tau\Omega})\|_{2}$ $\leq\frac{3\sigma_{m}(\Omega)}{4}\,\epsilon\,,$ (170)
$\|\Omega(\mathds{1})-\widehat{\Omega(\mathds{1})}\|_{2}$ $\leq\frac{\sqrt{3}\sigma_{m}(\Omega)}{2\sqrt{d_{B}}}\epsilon\,,$ (171)
@@
It is immediate to verify that the above inequalities are such that $\|\Omega-\hat{\Omega}\|_{2\rightarrow 2}\leq\|\Omega-\hat{\Omega}\|_{2}\leq\frac{\sigma_{m}(\Omega)^{2}}{3}\leq\frac{\sigma_{m}(\Omega)}{3}$, because $\sigma_{m}(\Omega)\leq\|\Omega\|_{2}\leq 1$. We can thus insert the bounds from Proposition C.4 in the bound of Theorem C.3, and use $(1+a/t)^{t}\leq 1+2a$ for $0\leq a\leq 1$. ∎
The proof of the main theorem is analogous to the one of Theorem C.6. Note that the only difference is the dependence of the Hilbert-Schmidt error $\epsilon_{HS}$ on the parameters of the model, and see Definition 2.2 for the definiton of the class $\mathcal{S}_{q}(m,s,\eta)$.
Theorem C.9.
Let $\tilde{\omega}\in\mathcal{S}_{q}(d_{\mathcal{B}},s,\eta)$. Given $\hat{\omega}_{s},\hat{\omega}_{2s},\hat{\omega}_{2s+1}$ estimates of ${\omega}_{s},{\omega}_{2s},{\omega}_{2s+1}$ respectively, such that $D_{HS}(\hat{\omega}_{s},{\omega}_{s})$, $D_{HS}(\hat{\omega}_{2s},{\omega}_{2s})$, $D_{HS}(\hat{\omega}_{2s+1},{\omega}_{2s+1})$ are smaller than $\frac{\epsilon\eta^{3}}{20td_{\mathcal{B}}\sqrt{d_{\mathcal{A}}}}$. Then, $\hat{\omega}_{t}$ constructed from the output of Algorithm 1 satisfies
$\frac{\|\hat{\omega}_{t}-\omega_{t}\|_{1}}{2}\leq\epsilon$ (175)
Appendix D Non-translation-invariant FCS on a finite chain